Are these calculations correct to find the ROI, RER, Real Rate of Return and After-tax Rate of Return of an investment ?Over 3 years, an investment of $2500 increases to $3127. Last 36 months (3 years) inflation rate is 6.7 % (1.9+2.3+2.5) The tt.
Answer #1

ROI=Return on Investment =(Return/Investment )*100%=(3127-2500)/2500*100%=25.08%

ROR=Rate of Return =Annualized Return=r

(1+r)^3=(3127/2500)

1+r=(3127/2500)^(1/3)=1.077

ROR=r=1.077-1=0.077=7.7%

Real Rate of Return=Ir

Three Years annualized inflation=If

(1+If)^3=(1.019*1.023*1.025)=1.068498

1+If=1.068498^(1/3)=1.02233

Annualized Inflation Rate=If=0.02233

ROR=In=Nominal Rate of Return=0.077

(1+Ir)*(1+If)=(1+In)

1+Ir=(1+In)/(1+If)=1.077/1.02233=1.0535

Real Rate of Return =Ir=0.0535

Real Rate of Return =5.35%

Tax rate =22%=0.22

After Tax Real Rate of Return :

Taxes are paid on Nominal Return

After Tax Nominal Rate of Return =In*(1-Tax Rate)=0.077*(1-0.22)=0.06006

After Tax Real Rate on Return =(1.06006/1.02233)-1=0.0369

After Tax Real Rate of Return =3.69%